# Train Once, Deploy Anywhere: Realize Data-Efficient Dynamic Object Manipulation

### v1

@@ IV-D Hybrid Action Control
The control target of the end-effector tracking planner is to make the end-effector stably track the motion of the target object. In other words, the relative position offset between the end-effector and the target object should be maintained at the preset distances $(x^{p},y^{p},z^{p})$. In addition, to ensure the wrist RGB-D camera can clearly view the target object, the orientation of the end-effector should be kept at the preset Euler angles $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$. Summarily, the tracking action generated by the end-effector tracking planner comprises three parts, the tracking position $(x_{c}^{e}+x^{p},y_{c}^{e}+y^{p},z_{c}^{e}+z^{p})$, tracking orientation $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$, and tracking velocity $(v_{x},v_{y},v_{z})$.
Based on the tracking action, the end-effector first approaches the target object, as visualized in Moment 1 of Fig. 7. Then, at Moment 2, the end-effector achieves a stable tracking status. In this status, the moving target object appears stationary in the view of the wrist camera. After attaining this status, the learnable network (the recurrent probabilistic network) starts to predict manipulation actions for collision-rich interactions with the target object, as shown in Moment 3. Since the network generates manipulation actions based on the wrist-view 3D points $\{(x_{i}^{w},y_{i}^{w},z_{i}^{w})\}_{i=1}^{N^{w}}$, the original dynamic object manipulation problem is transformed into a static object manipulation problem in the view of the network. Hence, the motion pattern of the target object is excluded from the factors affecting observations, meaning $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$ is reduced. Besides, this design fixes the relative position between the end-effector and the target object in the beginning of manipulation action generation as $(x^{p},y^{p},z^{p})$, which further decreases $\mathcal{H}(\mathcal{O}_{\mathcal{C}})$.
The manipulation action predicted by the network includes the position offset $(x^{w},y^{w},z^{w})$ and orientation offset $(\theta^{w}_{x},\theta^{w}_{y},\theta^{w}_{z})$. The end-effector control target output by the hybrid action control block is the sum of the tracking action and manipulation action. Therefore, the end-effector control target consists of the position target $(x_{c}^{e}+x^{p}+x^{w},y_{c}^{e}+y^{p}+y^{w},z_{c}^{e}+z^{p}+z^{w})$, orientation target $(\theta^{p}_{x}+\theta^{w}_{x},\theta^{p}_{y}+\theta^{w}_{y},\theta^{p}_{z}+\theta^{w}_{z})$, and velocity target $(v_{x},v_{y},v_{z})$. Afterwards, the end-effector control target is converted into robot joint rotations based on inverse kinematics. This design enables a single GEM system to be applied to different robot embodiments.
Figure 8: The overall framework of the proposed recurrent probabilistic network, which takes the 3D points of concerned objects as input and includes a memory encoder to manage the historical observation memory.
IV-E Recurrent Probabilistic Network
Section IV-D has described how classic control and learnable network control are combined in the hybrid motion control block. This part further elaborates the structure details of the recurrent probabilistic network used in learnable network control. The overall framework of this network is illustrated in Fig. 8. To concentrate on the geometry of objects rather than appearance, the recurrent probabilistic network takes the wrist view 3D points produced by the object-centric geometry perception block as the main input. For the 3D points perceived at timestamp $t$, it first employs a point feature extractor to encode the 3D points into semantic-rich feature tokens $F_{t}\in\mathbb{R}^{l_{p}\times c}$, where $l_{p}$ denotes the token number and $c$ is the token length. In this work, we model the point feature extractor with Sonata [25], a recent promising point network. However, the input of Sonata should be colored points, while ours are uncolored points appended with instance IDs. To bridge this gap, we feed the instance IDs into Sonata as surrogate color information. Therefore, the points of various object instances in Fig. 8 are visualized with different colors.
@@ IV-E Recurrent Probabilistic Network
The memory encoder shown in Fig. 8 is to fuse $F_{t}$ and the memory feature at timestamp t1t-1, denoted as $M_{t-1}\in\mathbb{R}^{l_{m}\times c}$, into the new memory feature $M_{t}\in\mathbb{R}^{l_{m}\times c}$. To this end, the memory encoder first employs a Transformer decoder [59] to conduct feature interaction between $M_{t-1}$ and $F_{t}$. Specifically, $M_{t-1}$ is used as the query embedding, and $F_{t}$ is the key and value embeddings. The feature output by this Transformer decoder is transformed into the forget score $G_{t}\in\mathbb{R}^{l_{m}\times 1}$ and hidden feature $H_{t}\in\mathbb{R}^{l_{m}\times c}$ by two multilayer perceptron layers (MLPs). Afterwards, in the Mix module visualized in Fig. 8, $M_{t}$ is calculated following:
@@
where $\sigma(\cdot)$ represents the Sigmoid function. At the beginning of manipulation, the memory feature is initialized as a zero embedding. Taking these designs, the historical memory is updated in an auto-regressive style, and $\mathcal{H}(\mathcal{A}_{\mathcal{C}}|\mathcal{O}_{\mathcal{C}})$ is decreased.
Then, $M_{t}$, $F_{t}$, and a skill type embedding $S_{t}\in\mathbb{R}^{1\times c}$ are concatenated and input to three Transformer decoders. $S_{t}$ is a learnable token specifying which skill to perform, and the supported skills include Pick, Place, Rotate, and Insert. The feature interaction process in decoders is similar to the implementation in ACT [8], where a chunk of $N_{a}$ action tokens is utilized to conduct feature interaction with the input feature. Then, the action tokens are decoded as the action probability distributions of future $N_{a}$ timestamps and a binary finish flag. As described in Section III-E, the action probability distributions are represented as the means, variances, and component weights of Gaussian mixture distribution. The finish flag indicates whether the execution of the manipulation skill has been completed. If completed, the network stops generating actions until the next time of performing a skill. In training, the network is updated following Eq. (3) as described in Section III-E. During testing, the action with the maximum probability is selected for execution.
Figure 9: This figure illustrates our simulated dynamic object manipulation environment. (a) An overview of the environment, where diverse objects are moving on a conveyor belt. (b) We collect 3D assets in 30 representative geometric structures to collect demonstrations. (c) The Pick skill. (d) The Put skill. (e) The Rotate skill. (f) The Insert skill.
IV-F Efficient Demonstration Collection
Combining the aforementioned techniques, we derive a low-entropy manipulation system, which exhibits stronger generalization with less demonstration data. In this part, we introduce how the demonstration data is efficiently collected to support various manipulation skills within a single system.
Demonstration data can be collected in both simulated environments and the real world. Data collection in simulated environments is usually more efficient, but the visual fidelity of existing simulators is still limited. This limitation restricts existing simulated data generation. Thankfully, our network takes uncolored 3D points as input, avoiding the visual fidelity problem. Therefore, our method can generalize to real-world applications even though only using simulated data.
Nevertheless, there is no public simulated environment for dynamic object manipulation. To bridge this gap, we build the first such environment based on Isaac Gym [32]. As shown in Fig. 9 (a), the environment layout is similar to the real platform in Fig. 5. As mentioned in Section IV-C, geometric structure information of objects is essential for successful manipulation. To enable our method to generalize to objects in diverse geometries, we gather 3D assets of 30 representative geometric structures from the GSO dataset [60] to produce demonstration data. These 3D assets are visualized in Fig. 9 (b).
This work supports the four common skills including Pick, Place, Rotate, and Insert. These four skills are shown in Fig. 9 (c)-(f), respectively. Specifically, Pick requires the robot to pick up the moving object from the conveyor belt. Place needs the robot to put an object into a moving container on the conveyor belt. Rotate demands the robot to rotate the moving object to a target orientation. The target orientation in this work is prompted to the network by providing a reference observation, which is the point cloud of the manipulated object in the goal 3D orientation from the perspective of the network. Insert involves inserting an object into a container with a precisely sized opening, ensuring the object is aligned at the specified pose. The Insert task illustrated in Fig. 9 (f) only allows two millimeters of position error.
During gathering demonstrations, all the 30 assets can be used to collect Pick data. For Put, the 12 small-sized assets highlighted by a green box in Fig. 9 (b) are put into the 5 huge-sized containers marked by a yellow box. In Rotate, the 6 assets lacking rotational symmetry are utilized to collect data. For Insert, due to the size matching requirement between the insert object and the container, we only utilize the assets shown in Fig. 9 (f). Notably, the 3D assets are colored with random pure colors instead of being rendered with more realistic appearances. This is because our method takes uncolored point clouds as input, so there is no need to spend computational resources on rendering appearances.
Among these four skills, Pick requires human knowledge to decide reasonable grasp positions. For example, in Fig. 9 (c), the robot should grasp the hammer handle rather than the hammer head. Due to this requirement, the Pick data should be collected through human teleoperation. Specifically, we use our designed teleoperation handle shown in Fig. 5 (b) to collect 20 demonstrations for every 3D asset. In this way, 600 Pick demonstrations are obtained. Different from Pick, the demonstration data of the other three skills can be generated automatically through preprogrammed manipulation procedures. This is because we have the true state information of all objects in simulated environments. Hence, we build an automatic pipeline to generate data for these three skills without human intervention. A total of 2000 demonstrations are derived for each of these three skills.
@@ V-A1 Implementation Details
We train the recurrent probabilistic network using the 6,600 demonstrations collected in Section IV-F for 10K optimization iterations. The training takes about 14 hours on a single RTX4090 GPU. The optimizer is AdamW [61] and the learning rate is 1e41e-4. The training batch size is 32. The image resolutions of the two RGB-D cameras are both $640\times 480$. The action chunk size $N_{a}$, memory token number $l_{m}$, and feature channel number $c$ are set to 20, 200, and 512, respectively. GEM controls the robot at a frequency of 20 HZ.
V-A2 Testing Protocols
We test GEM in both the simulated environment in Fig. 9 and the real platform in Fig. 5. The primary evaluation metric is the task completion success rate. In Pick, the objects in various geometric structures are successively conveyed on the conveyor belt. The success condition for Pick is the gripper successfully picks the moving object up. In Put and Insert, an object and a container are moving on a conveyor belt. Success is defined as the robot arm placing the object inside the container. For Rotate, the robot hand grasps and rotates the object on the conveyor belt to the target orientation. Success is achieved if the yaw, pitch, and roll errors are all within 5 degrees. The default moving speed of the conveyor belt is 0.1 m/s. To mitigate the influence of random errors, we conduct each experiment for 100 times to compute the average success rate.
Figure 10: This figure illustrates the 14 selected objects in diverse geometric structures for out-of-domain manipulation evaluation. All the assets are used for Pick. In Put, the small-sized assets highlighted in a green box are required to be put in the containers marked in yellow. The ten 3D assets lack of rotational symmetry are employed in Rotate.
@@ V-A2 Testing Protocols
Figure 11: The 18 objects in diverse geometries that are used to conduct the real-world manipulation experiments. The left 9 objects can be used for collect demonstrations and are used in in-domain evaluation. The right 9 objects are for out-of-domain validation. All the objects can be used in Pick. For Put, one of the small-sized objects highlighted in the green box is put in one of the containers marked in the yellow box. The USB adapter and spoon are employed in Rotate. The objects in the orange box are for Insert, which also only allows two millimeters of manipulation error.

### v2

@@ IV-C Manipulation Acition Decomposition
The target of tracking actions is to make the end-effector stably track the motion of the target object. In other words, the relative position offset between the end-effector and the target object should be maintained at the preset distances $(x^{p},y^{p},z^{p})$. In addition, to ensure the wrist RGB-D camera can clearly view the target object, the orientation of the end-effector should be kept at the preset Euler angles $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$. Therefore, to maintain the synchronous motion of the robot end-effector and the moving target object in a relatively stationary state, the tracking action should include the tracking position $(x^{c}+x^{p},y^{c}+y^{p},z^{c}+z^{p})$, tracking orientation $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$, and tracking velocity $(v_{x},v_{y},v_{z})$.
By solely executing the tracking action as the manipulation action, the end-effector first approaches the target object, as visualized at Timestamp 1 of Fig. 5. Then, at Timestamp 2, the end-effector achieves a stable tracking status. In this status, the moving target object appears stationary in the view of the wrist camera. Upon attaining this stable status, as illustrated at Timestamp 3 of Fig. 5, the interaction action predicted by the GEM network begins to be sequentially added to the tracking action, which then serves as the composite manipulation action. The interaction action includes the position offset $(x^{n},y^{n},z^{n})$ and pose offset $(\theta^{n}_{x},\theta^{n}_{y},\theta^{n}_{z})$. In this control strategy, the tracking action alone ensures that the target object is in a relatively stationary state with respect to the robot end-effector. This means when the GEM network predicts interaction actions, it behaves as though it were manipulating a stationary object. In this way, the influence of various conveyor speeds on generating interaction actions is explicitly decoupled and removed. Therefore, this action decomposition strategy enables the GEM network trained using static object manipulation demonstrations to operate objects moving at diverse speeds on conveyor systems.
Summarily, before the end-effector achieves the stable tracking status, the manipulation action is identical to the tracking action computed from the point segmentation result, and it comprises the position goal $(x^{c}+x^{p},y^{c}+y^{p},z^{c}+z^{p})$, orientation goal $(\theta^{p}_{x},\theta^{p}_{y},\theta^{p}_{z})$, and velocity goal $(v_{x},v_{y},v_{z})$. After stable tracking is attained, the manipulation action is the sum of the tracking action and interaction action, including the position goal $(x^{n}+x^{c}+x^{p},y^{n}+y^{c}+y^{p},z^{n}+z^{c}+z^{p})$, orientation goal $(\theta^{n}_{x}+\theta^{p}_{x},\theta^{n}_{y}+\theta^{p}_{y},\theta^{n}_{z}+\theta^{p}_{z})$, and velocity goal $(v_{x},v_{y},v_{z})$. The manipulation action is transformed into the control signal of robot joints and end-effector based on inverse kinematics and PID [2]. Notably, the status flag predicted by the policy indicates whether the manipulation of an object has been completed. Once completed, the interaction action is removed from the manipulation action, and the tracking action switches to the next target object.
IV-D Network Implementation
Figure 6: The overall structure of the GEM network, which consists of four point encoders, three Transformer decoders, and a memory module. The memory module updates the memory representation in a recurrent manner. The network outputs include the GMM-based action distribution, point segmentation mask, and a status flag that marks the completion of manipulating one object.
The overall framework of the GEM network is illustrated in Fig. 6. During training, the network input is the 3D points perturbed by the appearance noise annealing strategy. In inference, the input is the points without noise. For the 3D points perceived at timestamp $t$, the network encodes the points into a semantic-rich feature embedding $F_{t}\in\mathbb{R}^{l_{p}\times c}$ using four point encoders [52], where $l_{p}$ denotes the token number and $c$ is the token length. We initialize the parameters of these encoders with Sonata [51]. Each token in $F_{t}$ corresponds to a voxel in the 3D physical space. In the point segmentation branch, a multilayer perceptron (MLP) layer is applied to transform $F_{t}$ into the segmentation mask by performing binary classification. The segmentation branch optimization objective is defined as a binary cross-entropy loss [26].
@@ IV-D Network Implementation
Subsequently, the memory module shown in Fig. 6 fuses $F_{t}$ and the memory feature at timestamp t1t-1, denoted as $M_{t-1}\in\mathbb{R}^{l_{m}\times c}$, into the new memory feature $M_{t}\in\mathbb{R}^{l_{m}\times c}$. During this process, the memory encoder first employs a Transformer decoder [5] to conduct feature interaction between $M_{t-1}$ and $F_{t}$. Specifically, $M_{t-1}$ serves as the query embedding, and $F_{t}$ is the key and value embeddings. The feature output by this Transformer decoder is transformed into the forget score $G_{t}\in\mathbb{R}^{l_{m}\times 1}$ and hidden feature $H_{t}\in\mathbb{R}^{l_{m}\times c}$ by two MLP layers. Afterwards, $M_{t}$ is calculated following:
@@
where $\sigma(\cdot)$ represents the Sigmoid function. At the beginning of manipulation, the memory feature is initialized as a zero embedding. Incorporating historical memory into the network observation is beneficial for enabling GEM to accurately assess the current manipulation progress, thereby reducing the ambiguity in network optimization.
Then, $F_{t}$ and $M_{t}$ are concatenated and input to three Transformer decoders to conduct feature interaction with $N_{a}$ action tokens. The interaction process in decoders is similar to the implementation of ACT [60], where the $N_{a}$ action tokens correspond to the $N_{a}$ actions scheduled for the subsequent $N_{a}$ timestamps. After the interaction in these decoders, we transform these action tokens into the probabilistic distributions of actions. This operation is different from recent works that commonly decode action tokens into deterministic action values [24]. We choose to predict action distributions because this design better captures the multi-modal nature of embodied actions, thereby alleviating the optimization conflict in training caused by decision-making ambiguity.
For each action token, we decode it into an action distribution represented using GMM with $K$ components. Specifically, an MLP layer predicts the means, variances, and weights of these $K$ Gaussian components based on each action token. For the $t_{\rm th}$ ($t\in[1,N_{a}]$) action token, denote the mean, variance, and weight of the $k_{\rm th}$ component as $\mu_{t,k}$, $\sigma_{t,k}^{2}$, and $\alpha_{t,k}$, respectively. The distribution of the $t_{\rm th}$ scheduled action $p(a_{t})$ can be formulated as:
$p(a_{t})=\sum\limits_{k=1}^{K}\alpha_{t,k}\mathcal{N}(\mu_{t,k},\sigma^{2}_{t,k}).$ (8)
The component weights $\{\alpha_{t,k}\}_{k=1}^{K}$ satisfy $\sum\limits_{k=1}^{K}\alpha_{t,k}=1$ and $\alpha_{t,k}\geq 0$ for all $k\in\{1,2,\cdots,K\}$. In training, assuming the corresponding expert actions of the $N_{a}$ timestamps are $\{\tilde{a}_{t}\}_{t=1}^{N_{a}}$. According to Eq. (3), the supervision loss for action prediction is defined as:
$L_{a}=-\frac{1}{N_{a}}\sum\limits_{t=1}^{N_{a}}\log(\sum\limits_{k=1}^{K}\frac{\alpha_{t,k}}{\sqrt{2\pi\sigma_{t,k}^{2}}}\exp(-\frac{(\tilde{a}_{t}-\mu_{t,k})^{2}}{2\sigma_{t,k}^{2}})).$ (9)
During inference, the highest-probability action at each of the $N_{a}$ future timestamps is selected, forming an action chunk that is executed in a single rollout. Moreover, the GEM network predicts a Boolean status flag that indicates whether the manipulation of the current object is completed. As described in Section IV-C, this flag controls when the policy should switch to the next target object. In training, the status flag prediction is supervised with a binary cross-entropy loss.
IV-E Data Generation in Simulation
[…]
Combining the aforementioned techniques, we derive an efficient policy that can realize sim-to-real generalization. In this subsection, we explain our built simulation environment for dynamic object manipulation on conveyor belts.
Thanks to the appearance noise annealing strategy, the GEM network primarily focuses on the geometric structure of objects, and its reliance on visual appearance is reduced. To ensure that the trained network can generalize to diverse unseen objects, we need to generate manipulation demonstrations using 3D assets of varied geometries. This is because a key advantage of our method is that while the texture appearances of objects are highly diverse and nearly impossible to exhaustively enumerate, the set of common geometries is relatively limited. Therefore, by collecting a set of 3D assets that covers common object geometries, our method can achieve promising generalization. This insight empowers us to avoid collecting a very large number of 3D assets, significantly reducing the manual effort required for 3D asset preparation. Based on this insight, we select 30 assets with representative geometries from the GSO dataset [10] to generate demonstration data, and these assets are depicted in Fig. 7 (a).
We adopt Isaac Gym [31] to build the simulation environment for dynamic object manipulation, as visualized in Fig. 7 (b). We do not employ more visually photorealistic simulators like Isaac Sim [34] because we find their simulation fidelity on object interaction dynamics is inadequate for automatic data generation. Hence, in this work, Isaac Sim is mainly utilized to evaluate the out-of-domain generalization capabilities of policies trained using the demonstrations collected in Isaac Gym. Since Isaac Gym does not support rendering complex appearance textures, we assign random solid colors to objects during data collection, as illustrated in Fig. 7 (b). By contrast, as depicted in Fig. 7 (a), objects exhibit realistic textures during the evaluation in Isaac Sim.
[…]
To enable automatic demonstration generation without human intervention, we employ CuRobo [45] to generate manipulation trajectories based on the true states of objects obtained from the simulator. We collect 2,000 demonstrations for each of the four skills, and every demonstration comprises about ten seconds of manipulation actions, robot proprioception information, and environment observations.
@@ V-A1 Implementation Details
As described in Section IV-E, we generate 2,000 demonstrations for each skill in Pick, Put, Rotate, and Insert. We obtain a total of 8,000 demonstrations for these four skills and utilize them to train networks. We optimize the GEM network for 10K optimization iterations. The training takes about 14 hours on two RTX4090 GPUs. The optimizer is AdamW [30] and the learning rate is 1e41e-4. The training batch size is 32. The image resolutions of the two RGB-D cameras are both $640\times 480$. The action chunk size $N_{a}$, memory token number $l_{m}$, and feature channel number $c$ are set to 20, 200, and 512, respectively. GEM controls the robot at a frequency of 20 HZ. Since objects are transported successively, this work does not consider scenarios where objects on the conveyor belt are cluttered together.
Figure 8: The real-robot experiment platform and remote teleoperation handle. Specifically, Sub-figure (a) illustrates the real-robot environment for studying dynamic object manipulation. GEM controls the robot arm to perform diverse tasks given the observations from the external and wrist RGB-D cameras. Sub-figure (b) shows our designed remote handle for teleoperation. In the experiments needing real-world tuning data, we utilize this remote handle to collect demonstrations efficiently.
@@
Figure 9: The 18 objects of diverse geometries selected to conduct real-world experiments. All the objects can be used in the evaluation of the skills Pick and Rotate. For Put, one of the small-sized objects highlighted in the green box is put in one of the containers in the yellow box. The objects in the orange box are for Insert, which only allows two millimeters of insertion error.
Figure 10: (a) We build a simulation environment in Isaac Sim to evaluate the out-of-domain generalization of policies trained using the demonstrations from Isaac Gym. (b) We select 20 additional assets to validate the manipulation performance of policies on unseen objects.
V-A2 Real-Robot Platform
To conduct real-robot experiments, we build a hardware platform as shown in Fig. 8. We employ this platform to evaluate the sim-to-real performance of policies trained using simulated demonstrations. In this platform, objects are driven by a conveyor belt at an adjustable velocity. A robot arm with a gripper is mounted on one side of this conveyor belt to manipulate the target object. Two RGB-D cameras are deployed to capture environment observations, one on the robot arm (the wrist camera) and the other on a fixed third-person perspective bracket (the external camera). GEM controls the robot arm to manipulate objects based on the observations from these two cameras.
As illustrated in Fig. 8 (b), a teleoperation handle is designed to teleoperate the robot arm in Fig. 8 (a). This handle is used in the experiments needing real-world tuning demonstrations. Specifically, we control the motion of the robot arm end-effector based on the relative pose change of the handle derived from the SLAM camera. The opening and closing of the gripper is regulated through the trigger pull degree. A battery is mounted on the handle to power all the devices in this handle. The control signals from the SLAM camera and trigger are wirelessly transmitted to the robot arm by the main control board, thereby enabling flexible remote teleoperation. The handle is not used when testing the method.
To evaluate the capabilities of GEM on manipulating objects in various geometries, we select 18 objects as shown in Fig. 9 to conduct experiments. All the objects can be used to validate the skills Pick and Rotate. For Put, the robot needs to put one of the small-sized objects highlighted in the green box into one of the containers marked in the yellow box. The objects in the orange box are for Insert, which are in various sizes and only allow two millimeters of insertion error.
V-A3 Simulation Platform
Although real-robot experiments directly reflect robot performance in the real world, it is inconvenient to adjust environment variables to conduct an ablation study. In contrast, simulation experiments allow controlling environmental variables precisely, so we also evaluate our method in simulation. As illustrated in Fig. 7 (a) and (b), we gather objects of diverse geometries and build a simulation environment in Isaac Gym to generate training demonstrations. We can also use this environment to validate the in-domain performances of trained policies.
To further test the out-of-domain generalization results, we develop a simulation environment in Isaac Sim, as depicted in Fig. 10 (a). We can find that the visual fidelity of the rendered environment in Isaac Sim is promising, ensuring the behaviors of the evaluated policies are similar to those in the real world. Notably, a problem with using Isaac Sim is the occasional occurrence of abnormal collision dynamics. For example, a gripper might lightly touch a cup, causing it to jump or move unrealistically. Due to this problem, we do not choose Isaac Sim to build the automatic demonstration generation pipeline. During evaluating policies in Isaac Sim, we manually inspect such simulation anomalies and exclude abnormal cases from the experimental statistics. In addition, as shown in Fig. 10 (b), we select 20 new objects to validate the manipulation success rates of policies on unseen objects.
V-A4 Testing Protocols
We analyze policies using both the aforementioned simulated environments and a real-world platform. The primary evaluation metric is the task completion success rate. In Pick, the objects in various geometric structures are successively conveyed on the conveyor belt. The success condition for Pick is the gripper successfully picks the moving object up. In Put and Insert, an object and a container are moving on a conveyor belt. Success is defined as the robot arm places the object inside the container. For Rotate, the robot hand grasps and rotates the object on the conveyor belt to the target orientation. Success is achieved if the yaw, pitch, and roll errors are all within 5 degrees. The default moving speed of the conveyor belt is 0.1 m/s. To mitigate the influence of random errors, we repeat each experiment for 100 times to compute the average success rate.
In simulation, we validate policies based on the in-domain and out-of-domain settings. In the in-domain setting, the policies are evaluated in Isaac Gym to manipulate the objects that have appeared in training demonstrations. By contrast, for the out-of-domain setting, the policies manipulate the unseen objects shown as Fig. 10 (b) in Isaac Sim environments. For real-robot experiments, we test policies with the tuned sim-to-real and zero-shot sim-to-real settings. In tuned sim-to-real, the policies pre-trained with simulated data are further fine-tuned with 10 real-world demonstrations before evaluation. Differently, the zero-shot sim-to-real setting requires the policies to operate a real robot without real-world data tuning.